Circular No. 5966

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994I IN NGC 5194 M. Richmond, Princeton University, reports on a possible progenitor

for SN 1994I: “Hubble Space Telescope PC images taken on 1992 July 14 show a stellar object at a position close to that of the supernova, α = 13^{h}27^{m}47^{s}.9, δ = +47^{o}26′59′′ (equinox 1950.0; based on the FITS header), in fair agreement with the positions reported on IAUC 5961 and 5963. Two images yield HST instrumental F555W magnitudes of ∼ 23.8. An HST image taken with the F664N filter on 1991 Dec. 20 (exposure not very long) shows no Hα emission near the position; a red progenitor is possibly favored.”

B. Schmidt, P. Challis, and R. Kirshner, Center for Astrophysics, report: “A spectrogram obtained with the F. L. Whipple Observatory 1.5-m telescope (+ FAST) on Apr. 5 by P. Berland confirms the conclusion by Filippenko et al. (IAUC 5964) that SN 1994I is not a type-II supernova, as had been reported earlier. The spectrum has strong lines at observed wavelengths 745, 611, 569, 492, 478, 434, and 377 nm, and no hydrogen is detected. M. Phillips, Cerro Tololo Interamerican Observatory, has pointed out that this object’s spectrum resembles the pre-maximum spectrum of the type-Ic object SN 1983V.”

R. A. Sramek and M. P. Rupen, National Radio Astronomy Observatory; S. D. Van Dyk, University of California at Berkeley; K. W. Weiler, Naval Research Laboratory; and N. Panagia, Space Telescope Science Institute, report: “New observations with the Very Large Array in its highestresolution (A) configuration have detected SN 1994I for the first time at 2 and 6 cm, and show that the 1.3- and 3.6-cm flux densities are rising rapidly. Following are listed the time of observation (UT), the frequency (GHz), the flux density (mJy), and the rms noise in image (mJy/beam): Apr. 4.11, 8.44, 0.91 ± 0.16, 0.054; 4.15, 14.94, 3.16 ± 0.52, 0.11; 4.16, 22.46, 4.64 ± 0.80, 0.42; 5.29, 1.43, <0.28 (3σ), 0.093; 5.32, 22.46, 13.1 ± 3.5, 0.26; 5.36, 14.94, 7.15 ± 1.05, 0.14; 5.38, 8.44, 1.66 ± 0.15, 0.044; 5.40, 4.86, 0.57 ± 0.08, 0.062. The uncertainties include both random and systematic (calibration) errors. The basic behavior—the flux densities rising rapidly, first at higher frequencies and later at lower frequencies—is consistent with the circumstellar interaction model and with the behavior of previously-known radio supernovae (as discussed, e.g., in Weiler et al. 1986, Ap.J. 301, 790).”

1994 April 6 Daniel W. E. Green